cathecat88:

John want to simplify the expression (5 + 3) (With the power of 2). As a first step, he writes 5( with a power of 2) + 3( With a power 2) will he get the correct value for the expression? If not, what should he do to simplify the expression?

memeXpapi:

k

xxDeppressionxx:

5 to the power of 2 plus 3 to the power of 2

xxDeppressionxx:

then 25 plus 9 = 34

Angle:

@xxDeppressionxx Your answer is incorrect. The question is asking: Is (5 + 3)^2 equal to (5^2) + (3^2) ?

Angle:

Using the order of operations: PEMDAS We do the stuff on the inside of the Parentheses first (5+3)^2 = (8)^2 Then 8^2 = 8 x 8 because that is how exponents work. Then the question is asking IS 8 x 8 equal to the 34 that xxDeppressionxx got before?